OpenStudy (anonymous):

Find the missing geometric sequence: 9/4, 3/4, 1/4, 1/12, ? A 1/36 B 1/20 C 1/6 D 1/3

hero (hero):

Hint: 9/4*1/3 = 3/4 3/4*1/3 = 1/4 1/4*1/3 = 1/12

OpenStudy (anonymous):

What does this mean?

hero (hero):

It means there is a pattern.